OpenStudy (angelwings996):

How do the graphs y = 1/x and y = 5 / (x + 6) compare ?

OpenStudy (angelwings996):

@mathstudent55

OpenStudy (mathstudent55):

y = 1/x has an asymptote at x = 0 y = 5/(x + 6) has an asymptote at x = -6 So the difference in denominator, from x to x + 6 causes a translation of -6 in x (6 units to the left).

OpenStudy (angelwings996):

Okay thank you so much!